# Carleman Estimates and Controllability of Forward Stochastic Parabolic Equations with General Dynamic Boundary Conditions

### v1

@@ 1 Introduction and Main Results
$\begin{aligned} \lambda^{3}\mu^{4}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{3}z^{2}dxdt+\lambda^{3}\mu^{3}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{3}z_{\Gamma}^{2}d\sigma dt=&-\mathbb{E}\iint_{Q}[\mathbbm{1}_{G_{0}}\widehat{u}z+F_{1}\widehat{y}-F\cdot\nabla\widehat{y}+Z\widehat{v}_{1}]dxdt\\ &-\mathbb{E}\iint_{\Sigma}[F_{2}\widehat{y}_{\Gamma}-F_{\Gamma}\cdot\nabla_{\Gamma}\widehat{y}_{\Gamma}+\widehat{Z}\widehat{v}_{2}]d\sigma dt.\end{aligned}$ (2.23)
$\begin{aligned} &\,\lambda^{3}\mu^{4}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{3}z^{2}dxdt+\lambda^{3}\mu^{3}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{3}z_{\Gamma}^{2}d\sigma dt\\ &\leq\rho\Bigg[\lambda^{-3}\mu^{-4}\mathbb{E}\iint_{Q}\theta^{-2}\varphi^{-3}\widehat{u}^{2}dxdt+\lambda^{-2}\mu^{-2}\mathbb{E}\iint_{Q}\theta^{-2}\varphi^{-2}\widehat{v}_{1}^{2}dxdt\\ &\hskip 22.76228pt+\lambda^{-2}\mu^{-2}\mathbb{E}\iint_{\Sigma}\theta^{-2}\varphi^{-2}\widehat{v}_{2}^{2}d\sigma dt+\mathbb{E}\iint_{Q}\theta^{-2}\widehat{y}^{2}dxdt+\mathbb{E}\iint_{\Sigma}\theta^{-2}\widehat{y}_{\Gamma}^{2}d\sigma dt\\ &\hskip 22.76228pt+\lambda^{-2}\mu^{-2}\mathbb{E}\iint_{Q}\theta^{-2}\varphi^{-2}|\nabla\widehat{y}|^{2}dxdt+\lambda^{-2}\mu^{-2}\mathbb{E}\iint_{\Sigma}\theta^{-2}\varphi^{-2}|\nabla_{\Gamma}\widehat{y}_{\Gamma}|^{2}d\sigma dt\Bigg]\\ &\quad+\frac{1}{4\rho}\Bigg[\lambda^{3}\mu^{4}\mathbb{E}\iint_{Q_{0}}\theta^{2}\varphi^{3}z^{2}dxdt+\lambda^{2}\mu^{2}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{2}Z^{2}dxdt\\ &\hskip 45.52458pt+\lambda^{2}\mu^{2}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{2}\widehat{Z}^{2}d\sigma dt+\mathbb{E}\iint_{Q}\theta^{2}F_{1}^{2}dxdt+\mathbb{E}\iint_{\Sigma}\theta^{2}F_{2}^{2}d\sigma dt\\ &\hskip 45.52458pt+\lambda^{2}\mu^{2}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{2}|F|^{2}dxdt+\lambda^{2}\mu^{2}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{2}|F_{\Gamma}|^{2}d\sigma dt\Bigg].\end{aligned}$ (2.24)
$\begin{aligned} &\,\lambda^{3}\mu^{4}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{3}z^{2}dxdt+\lambda^{3}\mu^{3}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{3}z_{\Gamma}^{2}d\sigma dt\\ &\leq C\Bigg[\lambda^{3}\mu^{4}\mathbb{E}\iint_{Q_{0}}\theta^{2}\varphi^{3}z^{2}dxdt+\lambda^{2}\mu^{2}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{2}Z^{2}dxdt\\ &\hskip 28.45274pt+\lambda^{2}\mu^{2}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{2}\widehat{Z}^{2}d\sigma dt+\mathbb{E}\iint_{Q}\theta^{2}F_{1}^{2}dxdt+\mathbb{E}\iint_{\Sigma}\theta^{2}F_{2}^{2}d\sigma dt\\ &\hskip 28.45274pt+\lambda^{2}\mu^{2}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{2}|F|^{2}dxdt+\lambda^{2}\mu^{2}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{2}|F_{\Gamma}|^{2}d\sigma dt\Bigg],\end{aligned}$ (2.25)
for all $\mu\geq C$ and $\lambda\geq C(e^{2\mu|\psi|_{\infty}}T+T^{2})$.
On the other hand, computing $d|\theta\varphi^{\frac{1}{2}}z|^{2}_{L^{2}(G)}$ by using Itô’s formula, integrating the equality with respect to $t\in(0,T)$ and taking the expectation on both sides, we obtain that
Similarly to (2.26), by computing $d|\theta\varphi^{\frac{1}{2}}z_{\Gamma}|^{2}_{L^{2}(\Gamma)}$, integrating the result on $(0,T)$ and taking the expectation, we get that
Combining (2.26) and (2.27), it follows that
which implies that
It is easy to see that for large $\lambda\geq CT^{2}$, we have
$|(\theta^{2}\varphi)_{t}|\leq CT\lambda e^{2\mu|\psi|_{\infty}}\theta^{2}\varphi^{3}\quad\;\;\textnormal{and}\quad\;\;|\nabla(\theta^{2}\varphi)|\leq C\lambda\mu\theta^{2}\varphi^{2}.$ (2.30)
$\begin{aligned} &\,\mathbb{E}\iint_{Q}\theta^{2}\varphi|\nabla z|^{2}dxdt+\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi|\nabla_{\Gamma}z_{\Gamma}|^{2}d\sigma dt\\ &\leq C\lambda\mu\mathbb{E}\iint_{Q}\theta^{2}\varphi^{2}|z||\nabla z|dxdt+CT\lambda e^{2\mu|\psi|_{\infty}}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{3}z^{2}dxdt\\ &\quad\,+CT\lambda e^{2\mu|\psi|_{\infty}}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{3}z_{\Gamma}^{2}d\sigma dt+C\mathbb{E}\iint_{Q}\theta^{2}\varphi|z||F_{1}|\,dxdt\\ &\quad\,+C\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi|z_{\Gamma}||F_{2}|\,d\sigma dt+C\mathbb{E}\iint_{Q}\theta^{2}\varphi|F||\nabla z|dxdt\\ &\quad\,+C\lambda\mu\mathbb{E}\iint_{Q}\theta^{2}\varphi^{2}|z||F|dxdt+C\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi|F_{\Gamma}||\nabla_{\Gamma}z_{\Gamma}|d\sigma dt\\ &\quad\,+C\mathbb{E}\iint_{Q}\theta^{2}\varphi Z^{2}dxdt+C\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi\widehat{Z}^{2}d\sigma dt.\end{aligned}$ (2.31)
$\begin{aligned} &\,\mathbb{E}\iint_{Q}\theta^{2}\varphi|\nabla z|^{2}dxdt+\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi|\nabla_{\Gamma}z_{\Gamma}|^{2}d\sigma dt\\ &\leq\varepsilon\mathbb{E}\iint_{Q}\theta^{2}\varphi|\nabla z|^{2}dxdt+\varepsilon\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi|\nabla_{\Gamma}z_{\Gamma}|^{2}d\sigma dt+\frac{C}{\varepsilon}\lambda^{2}\mu^{2}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{3}z^{2}dxdt\\ &\quad\,+\frac{C}{\varepsilon}\mathbb{E}\iint_{Q}\theta^{2}\varphi|F|^{2}dxdt+\frac{C}{\varepsilon}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi|F_{\Gamma}|^{2}d\sigma dt+CT\lambda e^{2\mu|\psi|_{\infty}}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{3}z^{2}dxdt\\ &\quad\,+CT\lambda e^{2\mu|\psi|_{\infty}}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{3}z_{\Gamma}^{2}d\sigma dt+C\lambda^{-1}\mu^{-2}\mathbb{E}\iint_{Q}\theta^{2}F_{1}^{2}\,dxdt+C\lambda\mu^{2}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{2}z^{2}\,dxdt\\ &\quad\,+C\lambda^{-1}\mu^{-1}\mathbb{E}\iint_{\Sigma}\theta^{2}F_{2}^{2}\,d\sigma dt+C\lambda\mu\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{2}z_{\Gamma}^{2}\,d\sigma dt+C\lambda\mathbb{E}\iint_{Q}\theta^{2}\varphi^{2}|F|^{2}dxdt\\ &\quad\,+C\lambda\mu^{2}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{2}z^{2}dxdt+C\mathbb{E}\iint_{Q}\theta^{2}\varphi Z^{2}dxdt+C\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi\widehat{Z}^{2}d\sigma dt.\end{aligned}$ (2.32)
Choosing a small enough $\varepsilon$ in (2.32), and multiplying the obtained inequality by $\lambda\mu^{2}$, we deduce that
$\begin{aligned} &\,\lambda\mu^{2}\mathbb{E}\iint_{Q}\theta^{2}\varphi|\nabla z|^{2}dxdt+\lambda\mu^{2}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi|\nabla_{\Gamma}z_{\Gamma}|^{2}d\sigma dt\\ &\leq C\lambda^{3}\mu^{4}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{3}z^{2}dxdt+C\lambda\mu^{2}\mathbb{E}\iint_{Q}\theta^{2}\varphi|F|^{2}dxdt+C\lambda\mu^{2}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi|F_{\Gamma}|^{2}d\sigma dt\\ &\quad\,+CT\lambda^{2}\mu^{2}e^{2\mu|\psi|_{\infty}}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{3}z^{2}dxdt+CT\lambda^{2}\mu^{2}e^{2\mu|\psi|_{\infty}}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{3}z_{\Gamma}^{2}d\sigma dt\\ &\quad\,+C\lambda^{2}\mu^{4}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{2}z^{2}\,dxdt+C\mathbb{E}\iint_{Q}\theta^{2}F_{1}^{2}\,dxdt+C\mu\mathbb{E}\iint_{\Sigma}\theta^{2}F_{2}^{2}\,d\sigma dt\\ &\quad\,+C\lambda^{2}\mu^{3}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{2}z_{\Gamma}^{2}\,d\sigma dt+C\lambda^{2}\mu^{2}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{2}|F|^{2}dxdt+C\lambda^{2}\mu^{4}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{2}z^{2}dxdt\\ &\quad\,+C\lambda\mu^{2}\mathbb{E}\iint_{Q}\theta^{2}\varphi Z^{2}dxdt+C\lambda\mu^{2}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi\widehat{Z}^{2}d\sigma dt.\end{aligned}$ (2.33)
Choosing a sufficiently large $\lambda\geq C(e^{2\mu|\psi|_{\infty}}T+T^{2})$ in (2.33), we conclude that
@@
$\begin{aligned} &\,\lambda\mu^{2}\mathbb{E}\iint_{Q}\theta^{2}\varphi|\nabla z|^{2}dxdt+\lambda\mu^{2}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi|\nabla_{\Gamma}z_{\Gamma}|^{2}d\sigma dt\\ &\leq C\lambda^{3}\mu^{4}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{3}z^{2}dxdt+C\lambda^{3}\mu^{3}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{3}z_{\Gamma}^{2}\,d\sigma dt+C\lambda^{2}\mu^{2}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{2}|F|^{2}dxdt\\ &\quad\,+C\lambda^{2}\mu^{2}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{2}|F_{\Gamma}|^{2}d\sigma dt+C\mathbb{E}\iint_{Q}\theta^{2}F_{1}^{2}\,dxdt+C\mu\mathbb{E}\iint_{\Sigma}\theta^{2}F_{2}^{2}\,d\sigma dt\\ &\quad\,+C\lambda^{2}\mu^{2}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{2}Z^{2}dxdt+C\lambda^{2}\mu^{2}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{2}\widehat{Z}^{2}d\sigma dt.\end{aligned}$ (2.34)
Finally, by combining (2.34) and (2.25), we obtain the desired Carleman estimate (1.6). This completes the proof of Theorem 1.1. ∎
3 Observability Problems
This section is addressed to proving the appropriate unique continuation property and observability inequality for the following adjoint backward equation of (1.7) in the sense introduced in [27, Chap. 7, Sect. 7.3]:
such that
$\|(z,z_{\Gamma})\|_{L^{2}_{\mathcal{F}}(\Omega;C([0,T];\mathbb{L}^{2}))}+\|(z,z_{\Gamma})\|_{L^{2}_{\mathcal{F}}(0,T;\mathbb{H}^{1})}+\|(Z,\widehat{Z})\|_{L^{2}_{\mathcal{F}}(0,T;\mathbb{L}^{2})}\leq C\|(z_{T},z_{\Gamma,T})\|_{L^{2}_{\mathcal{F}_{T}}(\Omega;\mathbb{L}^{2})}.$ (3.2)
Corollary 3.1.

### v2

@@ 2 Preliminaries and Well-Posedness Results
$\begin{aligned} &\int_{G}(z_{T}-z(t))\zeta\,dx+\int_{\Gamma}(z_{\Gamma,T}-z_{\Gamma}(t))\zeta_{\Gamma}\,d\sigma\\ &=\int_{t}^{T}\int_{G}\mathcal{A}\nabla z\cdot\nabla\zeta\,dx\,ds+\int_{t}^{T}\int_{\Gamma}\mathcal{A}_{\Gamma}\nabla_{\Gamma}z_{\Gamma}\cdot\nabla_{\Gamma}\zeta_{\Gamma}\,d\sigma\,ds\\ &\quad-\int_{t}^{T}\int_{G}\mathcal{A}\nabla h\cdot\nabla\zeta\,dx\,ds-\int_{t}^{T}\int_{\Gamma}\mathcal{A}_{\Gamma}\nabla_{\Gamma}h_{\Gamma}\cdot\nabla_{\Gamma}\zeta_{\Gamma}\,d\sigma\,ds\\ &\quad+\int_{t}^{T}\int_{G}F_{1}\zeta\,dx\,ds+\int_{t}^{T}\int_{\Gamma}F_{2}\zeta_{\Gamma}\,d\sigma\,ds\\ &\quad+\int_{t}^{T}\int_{G}Z\zeta\,dx\,dW(s)+\int_{t}^{T}\int_{\Gamma}\widehat{Z}\zeta_{\Gamma}\,d\sigma\,dW(s),\quad\text{a.s.}\end{aligned}$ (2.8)
Next, applying (2.4) to the third and fourth terms on the right-hand side of (2.8), we obtain
$\begin{aligned} &\int_{G}(z_{T}-z(t))\zeta\,dx+\int_{\Gamma}(z_{\Gamma,T}-z_{\Gamma}(t))\zeta_{\Gamma}\,d\sigma\\ &=\int_{t}^{T}\int_{G}\mathcal{A}\nabla z\cdot\nabla\zeta\,dx\,ds+\int_{t}^{T}\int_{\Gamma}\mathcal{A}_{\Gamma}\nabla_{\Gamma}z_{\Gamma}\cdot\nabla_{\Gamma}\zeta_{\Gamma}\,d\sigma\,ds\\ &\quad+\int_{t}^{T}\int_{G}F_{1}\zeta\,dx\,ds+\int_{t}^{T}\int_{\Gamma}F_{2}\zeta_{\Gamma}\,d\sigma\,ds\\ &\quad-\int_{t}^{T}\int_{G}F\cdot\nabla\zeta\,dx\,ds-\int_{t}^{T}\int_{\Gamma}F_{\Gamma}\cdot\nabla_{\Gamma}\zeta_{\Gamma}\,d\sigma\,ds\\ &\quad+\int_{t}^{T}\int_{G}Z\zeta\,dx\,dW(s)+\int_{t}^{T}\int_{\Gamma}\widehat{Z}\zeta_{\Gamma}\,d\sigma\,dW(s),\quad\text{a.s.}\end{aligned}$ (2.9)
@@
Fix $t\in(0,T)$. Applying Itô’s formula from Lemma 2.2 to $\|(z,z_{\Gamma})\|^{2}_{\mathbb{L}^{2}}$ and integrating over $s\in(t,T)$ followed by taking expectations, we obtain